OpenStudy (mary1999):

The area of a rectangle is 2916.The ratio of the width to the length is 4 colon 9.The perimeter of the rectangle is

OpenStudy (erak):

Okay, so let's look at the ratio first. If we set a variable like "n" to be the amount of length and width. A = lw 2916 = (4n)(9n) Does making that equation make sense?

OpenStudy (mary1999):

yes

OpenStudy (erak):

So if you multiplied the two terms on the right side and isolate for n, what do you get?

OpenStudy (mary1999):

36n

OpenStudy (erak):

In this case, it would be 36n^2, because that gets multiplied too.

OpenStudy (mary1999):

ok

OpenStudy (erak):

So what you have now is 2916 = 36n^2. Do you know how to get n by itself now?

OpenStudy (mary1999):

do you divide 2916 by 36?

OpenStudy (erak):

Yup, so that gives you 81 = n^2 What next?

OpenStudy (mary1999):

not sure

OpenStudy (erak):

So right now, in order to get "n" by itself, you need to get rid of that ^2, or squared function. What's the opposite to squaring?

OpenStudy (mary1999):

taking the square root?

OpenStudy (erak):

Yeah, so you would do like that so:\[\sqrt{81} = \sqrt{n ^{2}}\]\[9 = n\] Does everything make sense up until this step?

OpenStudy (mary1999):

yes I forget that last step

OpenStudy (erak):

Alright, so now you know the value of "n", what did we say about the ratio of length to width and how they are related to n?

OpenStudy (mary1999):

2916=(4n)(9n)

OpenStudy (erak):

Yeah, so you know width = 4n, and length = 9n. Use the value of n we just got to find out the width and length

OpenStudy (mary1999):

so I would do 4*9 and 9*9

OpenStudy (mary1999):

which is 36 and 81then I would do 2(36) 2(81)

OpenStudy (mary1999):

so p=234

OpenStudy (erak):

Yeah, nice job

OpenStudy (mary1999):

Thank you!! I always forget the squaring part